OpenStudy (anonymous):

PROBABILITY QUESTION: 15% of children attending kindergarten have not had a measles vaccine. In an class of 20 students, what is the probability that at least one has NOT had a measles vaccine?

OpenStudy (anonymous):

is it 0.96 which can also b written as 96%

OpenStudy (anonymous):

if that the right answer i will show how i got it

OpenStudy (isaiah.feynman):

Show it

OpenStudy (anonymous):

yes please show steps, thats the right answer

OpenStudy (anonymous):

ok first i found out the probability of getting 0 that have not had a measle vaccine using nCr *q^n-r*P^r

OpenStudy (anonymous):

where q= failures and p=success

OpenStudy (isaiah.feynman):

Binomial distribution!

OpenStudy (anonymous):

so the percent of failures is 85% because 15% have had not had a measle vaccine

OpenStudy (anonymous):

but in the equation write them as 0.85 and 0.15

OpenStudy (anonymous):

0.85= failures and 0.15=success, n=20 r=0

OpenStudy (anonymous):

when u plug in everything and u get 0.038759531

OpenStudy (anonymous):

okay makes sense so far

OpenStudy (anonymous):

then u would use a equation i came up with which is No Restriction - What you Dont Want where NR=1 and WYDW = P(no one who has not had a measle vaccine)

OpenStudy (anonymous):

1-P(0)= 1-0.038759531=0.96

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

okay yes it does

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

no problem